OpenStudy (anonymous):

How does (255/408) x 100= 62.5?

OpenStudy (texaschic101):

255/408 * 100 = 0.625 * 100 = 62.5

OpenStudy (texaschic101):

I turned the fraction into a decimal

OpenStudy (anonymous):

How did u do that?

OpenStudy (anonymous):

?